# Helix: Distributed Serving of Large Language Models via Max-Flow on Heterogeneous GPUs

### v1

@@ 5.5 Cluster Heterogeneity
This section evaluates Helix, under increased cluster heterogeneity, with offline serving of LLaMA 70B in a cluster of 42 nodes and 7 node types in simulation. Fig. 9(e) shows the decode throughput comparison. In this cluster, V100, T4, and T4$\times$2 nodes can not form a separate pipeline by themselves. We report the throughput without these machines for separate pipelines baseline. We also try to build a mixed pipeline using those machines and report the number with the mixed pipeline as separate pipeline+. Results show that Helix achieves 1.38$\times$, 2.72$\times$, and 2.11$\times$ throughput compared with Swarm, separate pipelines and separate pipelines+.
Figure 11: Left: decode throughput of Helix, Swarm and Random scheduling in single and distributed cluster setup. Right: a case study that illustrates the congestion in Swarm and Random scheduling. The model placement used is found by Helix.
5.6 Model Placement Deep Dive
In this section, we analyze the impact of different model placement methods on the serving throughput of the cluster. We test with LLaMA 70B in both single and distributed cluster setups using offline serving. We compare Helix’s model placement planner with that of Swarm’s and Petals’ [4], which is a decentralized heterogeneous LLM inference system. We compare with Petals here because it only has a model placement strategy with no centralized scheduling method and thus not directly applicable in end-to-end evaluation. Swarm divides the model layers into equal length stages and assigns nodes to stages so that each stage has equal amount of compute. Petals, on the other hand, allows nodes to carry different number of layers. It targets on decentralized LLM serving, and thus decides model placement for each node sequentially and greedily selects a placement that covers layers with the least compute. To avoid the impact of different scheduling methods so as to isolate the effect of model placement, we use Helix’s request scheduler for all methods. Results in Fig. 10 left show that Helix achieves 1.23$\times$ and 2.10$\times$ decode throughput against Petals and Swarm respectively in the single cluster setup. This matches our observation in the prototype system. For distributed clusters setup, Helix’s relative throughput is 1.34$\times$ and 2.49$\times$ compared with Petals and Swarm respectively. We use a case study to explain why Helix’s model placement method achieves the best performance.
@@ 5.6 Model Placement Deep Dive
Fig. 10 right shows the GPU compute utilization rate of each method’s model placement for serving LLaMA 70B in the single cluster setup. The model placement found by Swarm has a bottleneck at the end of its pipeline, where 4 T4 nodes each serves 4 layers. This bottleneck causes GPU under-utilization for A100 and L4 nodes, significantly decreasing the serving throughput. For the model placement of Petals, there is no apparent bottlenecks. However, the placement still under-utilizes 8 T4 nodes and 1 L4 node, negatively affecting the serving throughput. For the placement found by Helix, almost all nodes are full-utilized. The efficient use of compute nodes enables Helix to outperform the baselines by up to 2.10$\times$.
5.7 Request Scheduling Deep Dive
In this section, we analyze the impact of different request scheduling methods on the serving throughput of the cluster. We test with LLaMA 70B in both single and distributed cluster setups using offline serving. We compare Helix’s request scheduler with (1) Swarm, which chooses next level nodes with frequency proportional to their throughput, and (2) random scheduling, which randomly chooses a next level node in scheduling. To eliminate the impact of model placement, all methods use the model placement plan found by Helix. Results in Fig. 11 left shows that for single cluster setup, using Helix’s scheduling increases serving throughput by 23%. For distributed cluster setup, Helix’s throughput increase is 12%. The improvements for single cluster setup match our observation in the prototype system. Moreover, runtime monitoring shows that Swarm and random scheduling have severe congestion. We will further illustrate this in a case study.
@@ 5.7 Request Scheduling Deep Dive
Fig. 11 right shows the model placement plan from Helix’s model placement planner for LLaMA 70B in the distributed clusters setup. The model Plan avoids using the slow inter-cluster network connections as much as possible, but there are still a few compute nodes connected with low-bandwidth connections. When using Swarm or random scheduling to schedule requests, we observe severe congestion on the three links marked as “congestion” in the figure – prompt phase requests queue up on those links for an average of 5s - 16s before they can be transmitted. We root-cause the nodes responsible for the congestion and mark them orange in the figure. Surprisingly, we find that one congestion is caused by bad scheduling from a node 3 nodes away. This verifies the necessity of a global scheduling method that can take both network and compute into account. We also observe similar congestion when running Swarm’s request scheduling on the model placement it finds for LLaMA 70B in the distributed cluster setup.
Table 4: Problem size with and without pruning. var means variables, and const means constraints.
Problem size Without pruning With pruning
24-node 876 var 1122 const 1376 var 1848 const
42-node 2144 var 2772 const 4004 var 5502 const
(a) Cluster pruning
@@
Figure 12: Ablation study.
5.8 Ablation Study on Optimization
This section performs ablation study for the two MILP optimizations introduced in Sec. 3.4. We use the LLaMA 70B settings in Sec. 5.4 and 5.5 to perform the ablation. We refer to them as 24-node and 42-node setting.
@@ 5.8 Ablation Study on Optimization
When cluster pruning is enabled, we prune network connections such that the average degree of each node is 12. This number of connections is sufficient for LLM inference systems as we discuss below. This removes 50% network links for 24-node setting and 72% network links for 42-node setting. Table. 4 shows that pruning reduces problem size by around 40% and 50% for the two settings respectively. We also compare the offline decode throughput of the best model placement found with and without cluster pruning. Results in Fig. 12(a) shows that Helix achieves 16% and 4% better performance when we enable pruning in the two settings respectively, for this problem instance. We note that the amount of speed-up achieved would vary depending on the specific instance of the MILP problem at hand. The reason is that in LLM serving, network connections used is very sparse – usually each node only communicates with a few other nodes. Also, there are many equivalent model placements that can achieve the same throughput. Pruning the cluster very likely keeps some of these placements still valid. It makes the search for these placements easier with limited optimization time, as the problem size (and solution space size) is reduced.
@@
We compare the performance of running Helix’s model placement planner starting from solutions of heuristic methods and from default values. Since the best model placement found is the same for both methods, we compare the wall clock time to find the placement. Fig. 12(b) shows that running MILP from heuristic solutions takes 43% and 8% more time in the two setups respectively, for this problem instance. We note that the amount of speed-up achieved would vary depending on the specific instance of the MILP problem at hand. The results demonstrate that start from heuristic solutions expedites model placement in Helix.
6 Related Work
@@
There are a large number of works for serving machine learning models, discussing aspects including system implementation [24, 23], model placement [18, 28, 34, 43], and request scheduling [9, 48, 34]. However, due to LLM’s unique auto-regressive execution paradigm, these approaches fail to efficiently serve LLMs. Instead, many recent LLM-specific systems tackle the unpredictable execution time and high memory consumption in LLM serving. Orca [46] proposed iteration level scheduling to release resources once a request is finished. vLLM [16] introduced PageAttention to further reduce the memory consumption of each request by allocating exact number of pages it requires. Speculative Inference [17, 22] applies a small model to predict multiple output tokens, and verify them in a single iteration. Splitwise [27] and DistServe [50] found that disaggregating the prompt and decode phase can improve the throughput, since the two phases have different workload characteristics. Sarathi [2] introduced chunked prefill, which allocates a budget to the prompt phase to make each microbatch’s workload balanced, minimizing pipeline bubble. All above works are orthogonal to our work and can be integrated into our system, since our focus is on the cluster heterogeneity.
ML Workloads on Heterogeneous Cluster
@@
Several methods explored the potential of utilizing heterogeneous GPUs for ML tasks. Some of them [26, 14] co-design the model partition and placement on a heterogeneous cluster but assume a uniform network bandwidth. Learninghome [32] and DeDLOC [7] studied the network-aware routing on a decentralized cluster but only considers either data or pipeline parallelism individually. SWARM [31], as discussed in Sec. 2.1, optimized the pipeline communication in a heterogeneous network. However, it schedules only by the next stage’s metadata, lacking a global view. There are also several efforts on using approximations to reduce the network communication [42] or synchronization [12]. Most of them focus on model training. In model inference, especially LLMs, serving with heterogeneous and geo-distributed GPUs is not well studied. SkyPilot [45] and Mélange [8] selects the best type of GPUs for a request, but each request is served by a single GPU type. Petals [4], as discussed in Sec. 2.1, studies a decentralized pipeline parallel setup. It designs a greedy model allocation and request scheduling for a dynamical device group, losing optimizing opportunities for a fixed device group.
7 Conclusion
@@
This paper presents Helix, the first high-throughput, low-latency LLM serving engine for heterogeneous GPU clusters. Helix formulates and solves layer placement and request scheduling as a Max-Flow problem. Compared to existing solutions, Helix achieves significant improvements in throughput and latency.

### v2

@@ 6.5. High GPU-Heterogeneity Cluster
This section evaluates Helix on a highly heterogeneous GPU cluster with 42 compute nodes and 7 types of GPUs. Fig. 8 shows the results. In this cluster, V100, T4, and T4$\times$2 nodes cannot form serving pipelines by themselves. We report the throughput without those machines for SP. We also try to build a mixed pipeline using those machines and report the number with the mixed pipeline as SP+. Compared with Swarm, SP and SP+, Helix achieves 1.37$\times$, 2.91$\times$, and 2.24$\times$ throughput for offline serving, and 1.48$\times$, 3.29$\times$ and 2.54$\times$ for online serving. Helix also reduces prompt latency by 17%, 1% and 7% respectively. The average decode latency of Helix is slightly higher, because the baselines under-utilize the slow GPUs and serve most requests with fast GPUs with large VRAM. This result indicates that Helix can achieve consistently high performance when there are many types of GPUs in the cluster.
(a) Decode throughput.
(b) Request scheduling case study.
Figure 10. (a) Comparing different request scheduling methods with offline serving of LLaMA 70B. (b) A case study that illustrates the congestion in Swarm and random scheduling when serving LLaMA 70B in geo-distributed clusters setup.
[…]
Figure for request scheduling.
6.6. Model Placement Deep Dive
In this section, we analyze the impact of different model placement methods on the serving throughput of the cluster. We evaluate Helix’s offline serving performance on both single and geo-distributed clusters, and compare Helix’s model placement method with those used in Swarm and Petals. Both methods perform model placement using throughput-based heuristics. To isolate the effect of model placement, we use Helix’s request scheduler for all methods. Fig. 9(a) shows the decode throughput. Compared with Petals and Swarm, Helix achieves 1.23$\times$ and 2.10$\times$ higher throughput on the single cluster, and 1.49$\times$ and 2.38$\times$ throughput on the geo-distributed clusters. We perform a case study on LLaMA 70B to demonstrate why Helix achieves the best performance.
@@ 6.6. Model Placement Deep Dive
Fig. 9(b) shows the model placement and GPU compute utilization for each method when serving LLaMA 70B on the single cluster. Swarm’s model placement introduces a bottleneck at the end of its pipeline, where 4 T4 nodes each serves 4 layers. This bottleneck causes GPU under-utilization on A100 and L4 nodes, significantly decreasing the serving throughput. For Petals’ model placement, 8 T4 nodes and 1 L4 node are under-utilized, which negatively affects the serving throughput. For Helix’s model placement, almost all nodes are fully-utilized. The efficient use of GPUs enables Helix to outperform Swarm and Petals by 2.10$\times$ and 1.23$\times$ respectively.
6.7. Request Scheduling Deep Dive
This section analyzes the impact of different request scheduling methods on the serving throughput of the cluster. We evaluate Helix’s offline serving performance of LLaMA 70B on both the single cluster and geo-distributed clusters. We compare Helix’s request scheduler with (1) Swarm, which schedules requests based on real-time throughput of each candidate node, and (2) random scheduling, which randomly chooses a candidate in scheduling. For the geo-distributed clusters setup, we further compare with Shortest Queue First scheduling (SQ), which always assigns requests to the node with shortest queue. To eliminate the impact of model placement, all methods use the model placement found by Helix. Fig. 10(a) shows the decode throughput. Compared with Swarm and random scheduling, Helix achieves 30% and 29% higher throughput on the single cluster, and 22% and 15% higher throughput on the geo-distributed clusters. Compared with Shortest Queue First scheduling, Helix achieves 19% higher decode throughput. Moreover, runtime monitoring shows that all three baseline scheduling methods introduce severe congestion. We illustrate this further in the case study.
@@ 6.7. Request Scheduling Deep Dive
Fig. 10(b) shows the model placement plan found by Helix for serving LLaMA 70B on the geo-distributed clusters. The plan avoids using slow inter-cluster network connections as much as possible, but a few compute nodes are still connected with low-bandwidth connections. When using Swarm or random scheduling to schedule requests, we observe severe congestion on the three links marked as “congestion” in the figure – prompt phase requests queue up on those links for an average of 5s - 16s before they can be transmitted. We root-cause the nodes responsible for the congestion and mark them orange in the figure. Surprisingly, we find that one congestion is caused by bad scheduling from a node 3 hops away. This verifies the necessity of a global scheduling method that can take both network and compute into account. We also observe similar congestion when serving LLaMA 70B with Swarm’s request scheduling on the model placement it finds for the geo-distributed clusters.
Table 8. Problem size with and without pruning. var means variables, and cstr means constraints.
Problem size With pruning Without pruning
24-node 876 var 1122 cstr 1376 var 1848 cstr
42-node 2144 var 2772 cstr 4004 var 5502 cstr
(a) Cluster pruning.
@@
Figure 11. Ablation study on MILP optimization.
[…]
Figure for ablation study.
6.8. Ablation Study on Optimization
This section performs an ablation study on the two MILP optimizations introduced in Sec. 4.5. We evaluate offline serving of LLaMA 70B on the geo-distributed and highly heterogeneous clusters, referred to as the 24-node and 42-node settings respectively.
@@ 6.8. Ablation Study on Optimization
When cluster pruning is enabled, we prune network connections such that the average degree of each node is 12, which is sufficient for LLM inference systems as we discuss below. Enabling cluster pruning removes 50% and 72% network connections for 24 and 42-node settings. Table. 8 shows that pruning reduces problem size by 36% and 46% for the two settings. Fig. 11(a) shows that Helix achieves 16% and 2% higher decode throughput when using the model placement found with cluster pruning. We note that the amount of speed-up achieved would vary depending on the specific instance of the MILP problem at hand. Pruning slow network connections does not harm throughput because network connections used in serving is very sparse – usually each node only communicates with a few other nodes. Also, there are many equivalent model placements that can achieve the same throughput. Pruning the cluster very likely keeps some of these placements still valid. It makes the search for these placements easier with limited optimization time, as the problem size (and solution space size) is reduced.
@@
We compare the performance of running Helix’s model placement planner starting from solutions of heuristic methods and from default values. Since the best model placements found are the same, we compare the wall clock time to find the placement. Fig. 11(b) shows that running MILP from heuristic solutions takes 43% and 8% less time for the 24- and 42-node setup. We note that the speed-up achieved would vary depending on the specific instance of the MILP problem at hand. The results show that starting from heuristic solutions accelerates model placement in Helix.
6.9. Model Placement Quality
Figure 12. Best solution and best upper bound found by the MILP solver relative to solving time. The red dotted line marks the optimal throughput for this cluster.
[…]
Figure for model placement quality.
This section evaluates the quality of model placements found by Helix relative to the MILP solving time. In this evaluation, we run Helix to find the optimal model placement for serving LLaMA 30B in a cluster with 4 L4 and 6 T4 machines. We record the best model placement as well as the best upper bound found by the MILP solver that we used (Gurobi) during the solving process. The best upper bound represents the best possible objective value that could be achieved for the MILP problem and will gradually become tighter as the solver explores more nodes and adds cutting planes. Fig. 12 shows the quality of the best model placement and the best upper bound relative to the solving time. Results show that Helix finds the optimal solution in less than 5 minutes, but it takes the solver more than one hour to enumerate all possible solutions and confirm the optimality. This indicates that we can early-stop the solving process, as high-quality solutions emerge in the early stages of computation.
7. Related Work
@@
There are a large number of works for serving machine learning models, discussing aspects including system implementation (olston2017tensorflow, ; nvidiaTriton, ), model placement (li2023alpaserve, ; pope2023efficiently, ; shen2019nexus, ; wu2023transparent, ), request scheduling (gujarati2020serving, ; zhang2019mark, ; shen2019nexus, ), and tail-latency mitigation (kosaian2019parity, ; kosaian2020learning, ; narra2020collage, ). However, due to LLM’s unique auto-regressive execution paradigm, these approaches fail to efficiently serve LLMs. Instead, many recent LLM-specific systems tackle the unpredictable execution time and high memory consumption in LLM serving. Orca (yu2022orca, ) proposed iteration level scheduling to release resources once a request is finished. vLLM (kwon2023vllm, ) introduced PageAttention to further reduce the memory consumption of each request by allocating exact number of pages it requires. Speculative Inference (leviathan2023fast, ; miao2023specinfer, ) applies a small model to predict multiple output tokens, and verify them in a single iteration. Splitwise (patel2023splitwise, ) and DistServe (zhong2024distserve, ) found that disaggregating the prompt and decode phase can improve the throughput, since the two phases have different workload characteristics. Sarathi (agrawal2023sarathi, ) introduced chunked prefill, which allocates a budget to the prompt phase to make each microbatch’s workload balanced, minimizing pipeline bubble. All above works are orthogonal to our work and can be integrated into our system, since our focus is on the cluster heterogeneity.
ML workloads on heterogeneous clusters.
@@
Several methods have utilized heterogeneous GPUs for ML tasks. Some of them (park2020hetpipe, ; jia2022whale, ) co-design the model partition and placement on a heterogeneous cluster but assume a uniform network bandwidth. Learninghome (ryabinin2020towards, ) and DeDLOC (diskin2021distributed, ) studied the network-aware routing on a decentralized cluster but only considers either data or pipeline parallelism individually. SWARM (ryabinin2023swarm, ), as discussed in Sec. 2.1, optimized the pipeline communication in a heterogeneous network. However, it schedules only by the next stage’s metadata, lacking a global view. There are also several efforts on using approximations to reduce network communication (wang2023cocktailsgd, ) or synchronization (hsieh2017gaia, ). Most of them focus on model training. In model inference, especially LLMs, serving with heterogeneous and geo-distributed GPUs is not well studied. SkyPilot (yang2023skypilot, ) and Mélange (griggs2024m, ) select the best type of GPUs for a request, but each request is served by a single GPU type. Petals (borzunov2022petals, ), as discussed in Sec. 2.1, studies a decentralized pipeline parallel setup. It designs a greedy model allocation and request scheduling for a dynamical device group, losing optimizing opportunities for a fixed device group. HexGen (jiang2023hexgen, ) is a concurrent work on LLM serving in heterogeneous clusters. It is based on fixed pipelines (see Sec. 5.1) and adopts heuristic-based methods to search for the optimal model placement. In comparison, our Max Flow formulation and per-request pipeline is more flexible than HexGen and our MILP formulation can guarantee the optimal solution.
Scheduling for Heterogeneous Resources
There exists extensive research on scheduling algorithms for heterogeneous resources. For example, energy-aware scheduling (linux_energy_sched, ) in the Linux kernel schedules tasks for heterogeneous CPU topologies. There are also several works (topcuoglu2002performance, ; radulescu2000fast, ; bajaj2004improving, ) that focus on general scheduling in heterogeneous clusters. Due to the auto-regressive nature of LLM inference, these methods cannot be directly used for heterogeneous LLM serving.
8. Conclusion
@@
This paper presents Helix, the first high-throughput, low-latency LLM serving engine for heterogeneous GPU clusters, with guaranteed optimal solution that maximizes throughput. Helix formulates and solves the model placement and request scheduling as a Max-Flow problem. Compared to existing solutions, Helix achieves significant improvements in throughput and latency.
Acknowledgment
We thank the anonymous reviewers and our shepherd Íñigo Goiri for their valuable feedback and constructive suggestions that helped improve this paper. We also express our gratitude to the Google Cloud Innovator program for providing the machines on Google Compute Engine for our experiments. This work was supported in part by a Sloan Foundation Fellowship and a VMware Systems Research Award. This work was also partially supported by the National Science Foundation under grant numbers CNS-2147909, CNS-2211882, and CNS-2239351, along with gift awards from Amazon and Meta.
Appendix A Artifact
A.1. Abstract
Our artifacts include Helix’s simulator and prototype system for distributed LLM serving in heterogeneous and geo-distributed clusters. The code implements key algorithms including the MaxFlow-based LLM serving formulation, MILP-based model placement planner, and per-request pipeline scheduler. We also provide comprehensive documentation and scripts for environment setup and system execution from scratch. The artifacts contain the identical simulator and prototype system used in Helix’s evaluation in the paper. The simulator can run on a single machine, while the prototype system requires a cluster deployment.
A.2. Artifact check-list (meta-information)
  • Algorithm: MaxFlow-based LLM serving formulation; MILP-based model placement planner; per-request pipeline scheduler
  • Compilation: The simulator is implemented in Python. The prototype system’s inter-node communication framework is written in C++ (recommended to compile with GCC 13.2) and includes an automated compilation script with CMake, while its remaining components are implemented in Python.
  • Model: We use LLaMa-2 70B as the test workload. The prototype system operates with dummy weights, requiring only the model architecture specification (which is provided in our code repository).
  • Data set: We use the Azure Conversation Dataset as our evaluation traces, with a pre-parsed version included in our code repository.
  • Run-time environment: For the simulator, we recommend using Python 3.10, while it can also support other recent Python versions. The simulator is not sensitive to OS versions. For the prototype system, we recommend using Ubuntu 24.04 LTS to setup the environment. We also recommend using conda to isolate the run-time environment for both systems. For detailed software dependencies, please refer to Sec. A.3.3.
  • Hardware: The simulator runs on a single machine without specific hardware requirements. The prototype system runs on a cluster of 24 machines with NVIDIA GPUs. Please refer to Sec. A.3.2 for detailed hardware requirements.
  • Metrics: We evaluate using the same metrics as presented in the paper: decoding throughput, decoding latency, and prompt processing latency.
  • Output: Results are logged to both terminal output and files. We include expected results in the code repository.
  • How much disk space required (approximately)?: The total size of all log files is around 300 MB.
  • How much time is needed to prepare workflow (approximately)?: The environment setup takes around 1 - 2 hours.
  • How much time is needed to complete experiments (approximately)?: Functionality evaluation takes around 2 hours. Reproducibility evaluation takes around 16 hours. (The part that needs the whole 24 machine cluster is around 4 hours. Namely, the sections that need the whole cluster are part of Sec 6.3, 6.6 and 6.7.)
  • Publicly available?: Yes. DOI: 10.5281/zenodo.14037926.
  • Code licenses (if publicly available)?: Apache 2.0
A.3. Description
A.3.1. How to access
The artifact is publicly available at https://github.com/Thesys-lab/Helix-ASPLOS25. It is also archived as 10.5281/zenodo.14037926. The file size is around 300 MB. Please refer to our Github repository for the latest version.
A.3.2. Hardware dependencies
The simulator runs on a single machine without specific hardware requirements. We recommend at least 32 GB of memory for simulating large clusters to prevent out-of-memory issues. The prototype system requires cluster deployment - our example configuration uses 24 machines, consisting of 4 machines with 1×A100-40GB, 8 machines with 1×L4, and 12 machines with 1×T4. We recommend to set up each machine with at least 16 CPU cores and 128 GB memory to avoid out-of-memory issues. The network connection between machines should ideally be at least 10 Gbps, with latency of approximately a few milliseconds. Usually machines in the same region from common cloud providers can meet the network requirements. For functionality test purposes, it is also possible to use machines from different regions.
A.3.3. Software dependencies
We recommend running the simulator with Python 3.10. It relies on networkx, matplotlib and gurobipy. The MILP-based model placement planner uses Gurobi as the MILP solver. Running the example in the code base does not require additional licenses. However, if you want to run model placement for larger clusters, it is necessary to acquire a Gurobi license that does not limit problem size. We recommend running the prototype system on Ubuntu 24.04 LTS and with Python 3.10. To build the inter-node communication framework, you need to install build-essential, cmake, libzmq, cppzmp and pybind11. To run the prototype system, you need to install CUDA 12.6 and vLLM 0.4.0.post1. For a step-by-step guide to setting up the environment and running the experiments, please refer to https://github.com/Thesys-lab/Helix-ASPLOS25/blob/master/readme.md.
A.3.4. Data sets
We use the Azure Conversation Dataset as our evaluation traces, with a pre-parsed version included in our code repository.
A.3.5. Models
We use LLaMa-2 70B as the test workload. The prototype system operates with dummy weights, requiring only the model architecture specification.
A.4. Detailed Steps for Reproducing Results
We provide an example to show the functionality of our system, please refer to https://github.com/Thesys-lab/Helix-ASPLOS25/blob/master/readme.md. We also provide another example to reproduce all result we got in the paper, please refer to https://github.com/Thesys-lab/Helix-ASPLOS25/blob/ master/artifact__\__evaluation/ae__\__readme.md